You purchase a $750,000 condo in Hoboken with a $550,000 permanent mortgage loan for 30 years. The annual interest rate on the loan is 4%. Approximately how much are your monthly payments?

a. $2,387

b. $2,625

c. $3,580

d. $19,295

e. $31,807
Answer #1

Rate per month =4%/12
Number of Periods =30*12 =360
PV of loan =550000
Monthly Payment =PV/((1-(1+r)^-n)/r) =550000/((1-(1+4%/12)^-360)/(4%/12)) =2625

Option b is correct option